OpenStudy (anonymous):

What is the volume of a sphere with a radius of 6 inches? a. 112π cubic inches b. 288π cubic inches c. 290π cubic inches d. 300π cubic inches

OpenStudy (mrnood):

all you need is th eformula for the volume which is V = 4/3 * pi * r^3 so put r= 6 into that formula and work it out

OpenStudy (mrnood):

you don't even need to put pi in the equation - because they have left it in the answer for you so the answer is pi * 6^3 *4 /3

OpenStudy (mrnood):

6^3 is the same as 6*6*6 no reason to wait for someone else to work it out for you....

OpenStudy (anonymous):

i still don't understand

OpenStudy (mrnood):

there is nothing to understand - and I am not going to give you the answer the formula to calculate the volume is \[V= \frac{ 4 }{ 3 } \times \pi \times r ^{3}\] the only variable in that equation is r - and you are told that r=6 so you need to calculate \[V = \frac{ 4 }{ 3 } \times \pi \times 6^{3}\] but just leave pi as a symbol (do not put 3.14.... in as a value) The answer will come out tobe one of the above options. You need to do some work to get the answer.

OpenStudy (anonymous):

904.779 cubic inches

OpenStudy (mrnood):

@Susan_Hester It is against the code of conduct to give direct answers. On top of that you have not read the question - and your answer is not given in the form of the answers in the post

OpenStudy (mrnood):

I am not stupid - I can do most of this maths blindfold and standing on my head. Please keep your insulting responses to yourself or you will be reported for abuse I repeat - do not give direct answers - the OP only learns that YOU know the answer from that. I also repeat that YOUR answer is not a suitable answer to the question, because you have not read it properly

OpenStudy (anonymous):

just BE QUIET like seriously @MrNood

OpenStudy (mrnood):

I suggest you read the code of conduct, and if you feel you can HELP someone with a question posted, then give them guidance towards solving the problem themselves.